# Cooperative Nonlinear Guidance Strategies for Guaranteed Pursuit-Evasion

### v2

@@ Abstract
This paper addresses the pursuit-evasion problem involving three agents– a purser, an evader, and a defender. We develop cooperative guidance laws for the evader-defender team that guarantee that the defender intercepts the pursuer before it reaches the vicinity of the evader. Unlike heuristic methods, optimal control, differential game formulation, and recently proposed time-constrained guidance techniques, we propose a geometric solution to safeguard the evader from the pursuer’s incoming threat. The proposed strategy is computationally efficient and expected to be scalable as the number of agents increases. Another alluring feature of the proposed strategy is that the evader-defender team does not require the knowledge of the pursuer’s strategy and that the pursuer’s interception is guaranteed from arbitrary initial engagement geometries. We further show that the necessary error variables for the evader-defender team vanish within a time that can be exactly prescribed prior to the three-body engagement. Finally, we demonstrate the efficacy of the proposed cooperative defense strategy via simulation in diverse engagement scenarios.
Index Terms:
@@
Pursuit-evasion, autonomy, aerospace, multiagent systems, aircraft defense.
I Introduction
@@
The pursuit-evasion problem has long been a subject of interest in game theory, robotics, aerospace, and control engineering. While the pursuit-evasion scenarios involving only two participants have been extensively explored, the extension to a three-body pursuit-evasion problem adds a layer of complexity and requires higher autonomy and strategic decision-making. Such engagements involve three agents, namely, a pursuer, a target, and a defender. The pursuer tries to capture the target, which in turn is assisted by a defender to avoid getting captured.
@@
Early research on the kinematics of three-body engagement could be found in [1]. In [1], a closed-form solution was derived for constant-bearing collision courses, while the work in [2] focused on determining the intercept point’s location in the evader-centered reference frame. For three-agent engagements, optimal control-based formulations with specific objectives, such as minimizing energy or cost, were employed in cooperative guidance strategies, as discussed in [3, 4, 5, 6, 7]. In [3], the authors presented a cooperative optimal guidance strategy integrated with a differential game formulation to maximize the separation between pursuer and evader. In [4], optimal cooperative pursuit-evasion strategies for the defender-evader team were proposed, considering arbitrary order-linearized dynamics for each agent. It was assumed that the pursuer’s guidance strategy was known in this case. The work in [5] introduced a multiple-model adaptive estimator approach for cooperative information-sharing between the evader and defender to estimate the likely linear guidance strategy of the pursuer. The work in [6] discussed three-layer cooperation between the defender and evader and explored information exchange between them, whereas that in [7] provided algebraic conditions under which the pursuer could capture the evader by bypassing the defender. Note that most of these strategies relied on linearized dynamics, simplifying guidance design but potentially limiting their applicability in diverse operating conditions and scenarios with significant heading errors.
@@
Guidance strategies developed in a nonlinear context can overcome these limitations and enhance performance, e.g., by relaxing small heading error assumptions and accounting for turn constraints. Notable works in this regard include [8, 9, 10, 11, 12]. In [8], the authors introduced a sliding mode control-based terminal intercept guidance and autopilot design for defenders to protect the evader from incoming pursuers. Another nonlinear guidance strategy employing sliding mode control was discussed in [9]. In [10], a nonlinear guidance strategy was explored for scenarios where multiple defenders simultaneously intercept the pursuer before it reaches the evader. In [11], nonlinear feedback laws were developed to guide the evader on a collision course with the pursuer as a decoy, allowing the defender to intercept the pursuer before it captures the evader. This approach also provided the defender with the flexibility to adopt either a defensive or aggressive stance based on mission requirements. Another nonlinear guidance strategy, based on relative line-of-sight error and time-to-go deviation between the evader and the defender, was presented in [12]. It is important to note that these guidance strategies relied on time-to-estimates, which may not always be available with the required precision, potentially affecting their effectiveness.
@@
Geometrical approaches have also found application in three-agent pursuit-evasion scenarios. For instance, in [13, 14], a method centered on line-of-sight guidance, a three-point guidance strategy, was explored. These three points were defined as the evader, the defender, and the pursuer. The approach demonstrated that if the defender remains aligned with the line-of-sight connecting the evader and the pursuer, the interception of the pursuer is assured before it nears the evader. In a related vein, a modified version of the command-to-line-of-sight guidance approach, incorporating optimal control theory and velocity error feedback, was introduced in [15]. Furthermore, in [16], the authors presented a guidance strategy based on a barrier Lyapunov function to protect the evader from a pursuer. It is worth noting that the use of the barrier function imposed restrictions on some engagement variables, such as the defender’s initial heading angle, thereby potentially limiting the target set in which the game could terminate. Motivated by these results, the focus of the current paper is analyzing and presenting a simple and intuitive geometry-based solution for guaranteed pursuit evasion. The merits of this work can be succinctly summarized as follows:
@@
  • Unlike LOS angle-based geometric guidance, wherein the defender has to strictly maintain a fixed angle of $\pi$ with respect to the pursuer-evader LOS, the proposed strategy is less stringent and only requires the said angle to be within a broad interval of $\left[{\pi}/{2},{3\pi}/{2}\right]$. This allows the defender to have more flexibility in desired angle selection depending on engagement scenarios.
@@
  • The proposed strategy has an intuitive appeal, is expected to be computationally efficient and sets itself apart from heuristic methodologies, optimal control strategies, and formulations rooted in differential games, where analytical solutions may cease to exist due to challenges associated with nonlinearity and the complexities inherent in non-convex optimization.
@@
Note that even if the pursuer is sufficiently close to the evader, the evader and defender will cooperatively maneuver such that the defender can intercept the pursuer. Of course, there could be a few cases where the pursuer is too close to the evader, while the defender is too far away from them. In such cases, it is apparent that the defender may not be able to defend the evader. However, it is also worthwhile to note that such cases are impractical to consider from a three-body aircraft defense scenario because if the defender is too far away, then the engagement, for all practical purposes, is only for two agents.
II Problem Formulation
We consider a cooperative defense problem involving nonholonomic agents, namely, a pursuer (P), an evader (E), and a defender (D). The pursuer aims to intercept the evader. In contrast, the defender’s objective is to neutralize the pursuer before it reaches the vicinity of the evader. Thus, the evader and the defender are allies, whereas the pursuer is the opposition. Such a scenario leads to an planar (2D) engagement, as shown in Figure 1. The agents evolve according to
$\dot{x}_{i}=v_{i}\cos\gamma_{i},~{}\dot{y}_{i}=v_{i}\sin\gamma_{i},~{}\dot{\gamma}_{i}=\dfrac{a_{i}}{v_{i}};\;\forall i={\mathrm{P},\mathrm{E},\mathrm{D}},$ (1)
where $[x_{i},y_{i}]^{\top}\in\mathbb{R}^{2}$, $v_{i}\in\mathbb{R}_{+}$, and $\gamma_{i}\in(-\pi,\pi]$ denote the position, speed, and the heading angle of the $i$th agent, respectively, whereas $a_{i}$ is its steering control (lateral acceleration), which is assumed to be bounded. Thus, $\left|a_{i}\right|\leq a_{i}^{\max}\in\mathbb{R}_{+}$. This is unlike previous studies where the agents have simple dynamics, and their instantaneous heading angles were used to control them [17, 18].
Figure 1: Three-body engagement.
Note that the considerations herein lead to nonlinear dynamics, including the vehicles’ turning constraints, and finding analytical solutions to guaranteed winning zones for either party is generally NP-hard. In this paper, the speeds and maneuverability of the agents are such that $v_{\mathrm{P}}\approx v_{\mathrm{D}}>v_{\mathrm{E}}$, and $a_{\mathrm{P}}^{\max}\approx a_{\mathrm{D}}^{\max}>a_{\mathrm{E}}^{\max}$, essentially implying that the pursuer and the defender are similar in capabilities with a speed advantage over the evader. As seen from Figure 1, the agents have relative separations, $r_{\ell}$, and the line-of-sight (LOS) angles between any two pair of agents is $\lambda_{\ell}$, where $\ell=\mathrm{EP},\mathrm{DE},$ and $\mathrm{DP}$. The engagement kinematics governing the relative motion between any two pairs can be expressed in polar coordinates as
@@ II Problem Formulation
$\dot{r}_{\mathrm{DE}}$ $=v_{\mathrm{E}}\cos\left(\gamma_{\mathrm{E}}-\lambda_{\mathrm{DE}}\right)-v_{\mathrm{D}}\cos\left(\gamma_{\mathrm{D}}-\lambda_{\mathrm{DE}}\right)$
$=v_{\mathrm{E}}\cos\delta_{\mathrm{ED}}-v_{\mathrm{D}}\cos\delta_{\mathrm{DE}}=v_{r_{\mathrm{DE}}},$ (2e)
$r_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}$ $=v_{\mathrm{E}}\sin\left(\gamma_{\mathrm{E}}-\lambda_{\mathrm{DE}}\right)-v_{\mathrm{D}}\sin\left(\gamma_{\mathrm{D}}-\lambda_{\mathrm{DE}}\right)$
$=v_{\mathrm{E}}\sin\delta_{\mathrm{ED}}-v_{\mathrm{D}}\sin\delta_{\mathrm{DE}}=v_{\lambda_{\mathrm{DE}}},$ (2f)
where $v_{r_{\mathrm{EP}}},v_{r_{\mathrm{DP}}},v_{r_{\mathrm{DE}}},v_{\lambda_{\mathrm{EP}}},v_{\lambda_{\mathrm{DP}}}$, and $v_{\lambda_{\mathrm{DE}}}$, represent the components of relative velocities of the relevant agents along and perpendicular to the corresponding LOS of their respective engagements. The quantities $\delta_{k}$ in (2) denote the corresponding lead angles and are defined as $\delta_{\mathrm{PE}}=\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}$, $\delta_{\mathrm{EP}}=\gamma_{\mathrm{E}}-\lambda_{\mathrm{EP}}$, $\delta_{\mathrm{PD}}=\gamma_{\mathrm{P}}-\lambda_{\mathrm{DP}}$, $\delta_{\mathrm{DP}}=\gamma_{\mathrm{D}}-\lambda_{\mathrm{DP}}$, $\delta_{\mathrm{ED}}=\gamma_{\mathrm{E}}-\lambda_{\mathrm{DE}}$, $\delta_{\mathrm{DE}}=\gamma_{\mathrm{D}}-\lambda_{\mathrm{DE}}.$ Note that (2a)–(2b) describe the equations of motion between the evader and the pursuer, while (2c)–(2d) represent the same for the defender-pursuer pair. On the other hand, the cooperative engagement between the evader-defender pair is described by (2e)–(2f).
@@ III Main Results
Let us define an angle $\chi$ such that it is the angle subtended by the evader-defender and the defender-pursuer LOS at any segment, which is measured positive in the counterclockwise sense. Referring to Figure 2, it is the $\angle$ E1D1P1 in $\triangle$ P1E1D1 and $\angle$ E0D0P0$=90^{\circ}$ in $\triangle$ P0E0D0. In the case of $\triangle$ P2E2D2, the fact that interior $\angle$ E2D2P2 is obtuse still holds but $\chi$ is exterior of this angle to respect the consistency of our definition. The following proposition provides a sufficient condition for the defender to capture the pursuer before the latter could intercept the evader.
@@
Consider the three-body engagement described using (2). If the defender maintains a positive closing speed with respect to the pursuer and attains an angle $\chi\in\left[\dfrac{\pi}{2},\dfrac{3\pi}{2}\right]$, then the pursuer’s capture is guaranteed before the evader could be captured.
@@
Referring to Figure 2, it is immediate that if $\chi\in\left(\dfrac{\pi}{2},{\pi}\right)$ represents a scenario depicted in $\triangle$ P1E1D1. It follows that the side E1P1 is the longest. Hence, E1P1$>$D1P1. If the evader-defender team cooperates such that the defender maintains a fixed $\chi\in\left(\dfrac{\pi}{2},{\pi}\right)$ while also ensuring a positive closing speed with respect to the pursuer ($v_{r_{\mathrm{DP}}}<0$), the $\triangle$ P1E1D1 will shrink in proportion and may change its orientation, thereby generating smaller similar triangles. Eventually, D1P1 will degenerate to zero (or $r_{\mathrm{D}_{1}\mathrm{P}_{1}}\to 0$) while E1P1 (or $r_{\mathrm{E}_{1}\mathrm{P}_{1}}$) will still be positive. By a similar argument, it readily follows that $r_{\mathrm{D}_{2}\mathrm{P}_{2}}\to 0$ before $r_{\mathrm{E}_{2}\mathrm{P}_{2}}$ if $\chi\in\left({\pi},\frac{3\pi}{2}\right)$. The cases of $\chi=\frac{\pi}{2},\pi$ are extremes but no different. When $\chi=\dfrac{\pi}{2}$ (as in $\triangle$ P0E0D0), the side E0P0 is the hypotenuse and still the longest, so same reasoning can be applied. Finally, when $\chi=\pi$, the defender is always directly between and on the pursuer-evader LOS, resulting in the pursuer’s capture by the defender prior to the interception of the evader by the pursuer. ∎
@@
$\chi=\pi+\lambda_{\mathrm{DP}}-\lambda_{\mathrm{DE}}$ (3)
for a general case. Toward this end, our goal is to ensure that the defender always attains a fixed angle, say, $\chi^{\star}\in\left[\frac{\pi}{2},\frac{3\pi}{2}\right]$. Thus, the control objective is to nullify the error variable,
$\beta=\chi-\chi^{\star}=\pi+\lambda_{\mathrm{DP}}-\lambda_{\mathrm{DE}}-\chi^{\star}.$ (4)
Additionally, we also desire that $\beta\to 0$ within a time that is independent and uniform with respect to the initial three-body engagement geometry for a guaranteed pursuit-evasion.
@@
The dynamics of the LOS angle in each pair of engagements have a relative degree of two with respect to the relevant steering controls in that engagement.
@@
On differentiating the LOS rate of evader-pursuer engagement given in (2b) with respect to time, one may obtain
Using the fact that $\dot{\delta}_{\mathrm{PE}}=\dot{\gamma}_{\mathrm{P}}-\dot{\lambda}_{\mathrm{EP}}$ and $\dot{\delta}_{\mathrm{EP}}=\dot{\gamma}_{\mathrm{E}}-\dot{\lambda}_{\mathrm{EP}}$, together with (1), the expression in (5) becomes
$r_{\mathrm{EP}}\ddot{\lambda}_{\mathrm{EP}}=$ $~{}-\dot{r}_{\mathrm{EP}}\dot{\lambda}_{\mathrm{EP}}+v_{\mathrm{P}}\cos\delta_{\mathrm{PE}}\left(\dfrac{a_{\mathrm{P}}}{v_{\mathrm{P}}}-\dot{\lambda}_{\mathrm{EP}}\right)$
$~{}-v_{\mathrm{E}}\cos\delta_{\mathrm{EP}}\left(\dfrac{a_{\mathrm{E}}}{v_{\mathrm{E}}}-\dot{\lambda}_{\mathrm{EP}}\right).$ (6)

### v3

# Senior $\memberSenior Member, IEEE$

@@ Abstract
This paper investigates a pursuit-evasion problem involving three agents: a pursuer, an evader, and a defender. Cooperative guidance laws are developed for the evader–defender team that guarantee interception of the pursuer by the defender before it reaches the vicinity of the evader. Unlike heuristic methods, optimal control, differential game formulation, and recently proposed time-constrained guidance techniques, a geometry-based solution is proposed to safeguard the evader from the pursuer’s incoming threat. The proposed strategy is computationally efficient and expected to be scalable as the number of agents increases. Another notable feature of the proposed strategy is that the evader–defender team does not require knowledge of the pursuer’s strategy, yet the pursuer’s interception is guaranteed for arbitrary initial engagement geometries. It is further shown that the relevant error variables for the evader–defender team (or individual) converge to zero at a prespecified finite time that can be exactly prescribed prior to the three-body engagement. Finally, the effectiveness of the proposed cooperative pursuit-evasion strategy is demonstrated through simulations across diverse engagement scenarios.
{IEEEkeywords}
@@
Pursuit-evasion, guidance, autonomy, aerospace, multi-agent systems, aircraft defense.
1 Introduction
@@
The pursuit-evasion problem has long been a subject of study in game theory, robotics, aerospace, and control engineering [1]. While the pursuit-evasion scenarios involving only two participants (see [2, 3] and references therein) have been extensively studied, the extension to a three-body pursuit-evasion problem adds a layer of complexity and requires higher autonomy and strategic decision-making. Such engagements involve three agents, namely, a pursuer, an evader, and a defender. The pursuer seeks to intercept the evader, while the defender assists the evader in avoiding interception.
@@
Early research on the kinematics of three-body engagement can be found in [4]. In [4], a closed-form solution was derived for constant-bearing collision courses, while the work in [5] focused on determining the intercept point’s location in the evader-centered reference frame. For three-agent engagements, optimal control-based formulations with specific objectives, such as minimizing energy or cost, were employed in cooperative guidance strategies, as discussed in [6, 7, 8, 9, 10]. In [6], the authors presented a cooperative optimal guidance strategy integrated with a differential game formulation to maximize the separation between pursuer and evader. In [7], optimal cooperative pursuit-evasion strategies for the defender-evader team were proposed, considering arbitrary order-linearized dynamics for each agent. It was assumed that the pursuer’s guidance strategy was known in this case. The work in [8] introduced a multiple-model adaptive estimator approach for cooperative information-sharing between the evader and defender to estimate the likely linear guidance strategy of the pursuer. The work in [9] discussed three-layer cooperation between the defender and evader and explored information exchange between them, whereas that in [10] provided algebraic conditions under which the pursuer could capture the evader by bypassing the defender. Note that most of these strategies relied on linearized dynamics, simplifying guidance design but potentially limiting their applicability in diverse operating conditions and scenarios with significant heading errors.
@@
Guidance strategies developed in a nonlinear context can overcome these limitations and enhance performance, e.g., by relaxing small heading error assumptions and accounting for turn constraints. Notable works in this regard include [11, 12, 13, 14, 15]. In [11], the authors introduced a sliding mode control-based terminal intercept guidance and autopilot design for defenders to protect the evader from incoming pursuers. Another nonlinear guidance strategy employing sliding mode control was discussed in [12]. In [13], a nonlinear guidance strategy was explored for scenarios where multiple defenders simultaneously intercept the pursuer before it reaches the evader. In [14], nonlinear feedback laws were developed to guide the evader on a collision course with the pursuer as a decoy, allowing the defender to intercept the pursuer before it captures the evader. This approach also provided the defender with the flexibility to adopt either a defensive or aggressive stance based on mission requirements. Another nonlinear guidance strategy, based on relative line-of-sight error and time-to-go deviation between the evader and the defender, was presented in [15]. It is important to note that these guidance strategies rely on time-to-estimates, which may not always be available with the required precision, potentially affecting their effectiveness.
@@
Geometrical approaches have also found application in three-agent pursuit-evasion scenarios. For instance, in [16, 17], a method centered on line-of-sight guidance, a three-point guidance strategy, was explored. These three points were defined as the evader, the defender, and the pursuer. The approach demonstrated that if the defender remains aligned with the line-of-sight connecting the evader and the pursuer, the interception of the pursuer is assured before it nears the evader. In a related vein, a modified version of the command-to-line-of-sight guidance approach, incorporating optimal control theory and velocity error feedback, was introduced in [18]. In [19], a two-stage guidance scheme based on the line-of-sight method for two-on-two engagements is presented. Furthermore, in [20], the authors presented a guidance strategy based on a barrier Lyapunov function to protect the evader from a pursuer. It is worth noting that the use of the barrier function imposed restrictions on some engagement variables, such as the defender’s initial heading angle, thereby potentially limiting the target set in which the game could terminate. Motivated by these results, the focus of the current paper is on analyzing and presenting a simple and intuitive geometry-based solution for guaranteed pursuit evasion. The merits of this work can be succinctly summarized as follows:
@@
  • Unlike LOS angle-based geometric guidance, wherein the defender has to strictly maintain a fixed angle of $\pi$ with respect to the pursuer-evader LOS, the proposed strategy is less stringent and only requires a said angle to be within a broad interval of $\left[{\pi}/{2},{3\pi}/{2}\right]$. This allows the defender to have more flexibility in desired angle selection depending on engagement scenarios.
  • We propose two levels of cooperation between the evader–defender team, which can be employed depending on the available communication between them and the required level of simplicity in implementation to safeguard the evader from the incoming pursuer.
@@
  • The proposed strategy has an intuitive appeal, is expected to be computationally efficient, and sets itself apart from heuristic methodologies, optimal control strategies, and formulations rooted in differential games, where analytical solutions may cease to exist due to challenges associated with nonlinearity and the absence of knowledge about the pursuer’s strategy.
@@
Note that even if the pursuer is sufficiently close to the evader, the evader and defender will cooperatively maneuver such that the defender can intercept the pursuer. Of course, there could be a few cases where the pursuer is too close to the evader, while the defender is too far away from them. In such pathological cases, it is apparent that the defender may not be able to defend the evader. However, it is also worthwhile to note that such cases are impractical to consider from a three-body aircraft defense scenario because if the defender is too far away, then the engagement, for all practical purposes, is only for two agents.
2 Problem Formulation
We consider a cooperative defense problem involving nonholonomic agents, namely, a pursuer (P), an evader (E), and a defender (D). The pursuer aims to intercept the evader. In contrast, the defender’s objective is to neutralize the pursuer before it reaches the vicinity of the evader. Thus, the evader and defender cooperate as allies, while the pursuer acts as the opponent. Such a scenario leads to a planar (2D) engagement, as shown in Figure 1. The agents evolve according to
$\dot{x}_{\ell}=v_{\ell}\cos\gamma_{\ell},~\dot{y}_{\ell}=v_{\ell}\sin\gamma_{\ell},~\dot{\gamma}_{\ell}=\dfrac{a_{\ell}}{v_{\ell}};\;\forall\ell\in\{\mathrm{P},\mathrm{E},\mathrm{D}\},$ (1)
where $x_{\ell},y_{\ell}]^{\top}\in\mathbb{R}^{2}$, $v_{\ell}\in\mathbb{R}_{+}$, and $\gamma_{\ell}\in(-\pi,\pi]$ denote the position, speed, and the heading angle of the $i$th agent, respectively, whereas $a_{\ell}$ is its steering control (lateral acceleration), which is assumed to be bounded. Thus, $\left|a_{\ell}\right|\leq a_{\ell}^{\max}\in\mathbb{R}_{+}$. This is unlike previous studies where the agents have simple dynamics, and their instantaneous heading angles were used to control them [21, 22].
Figure 1: Geometry of a three-body pursuit–evasion engagement.
In this paper, the speeds and maneuverability of the agents are such that $v_{\mathrm{P}}\approx v_{\mathrm{D}}>v_{\mathrm{E}}$, and $a_{\mathrm{P}}^{\max}\approx a_{\mathrm{D}}^{\max}>a_{\mathrm{E}}^{\max}$, essentially implying that the pursuer and the defender are similar in capabilities with a speed advantage over the evader. As seen from Figure 1, the agents have relative separations, $r_{\ell}$, and the line-of-sight (LOS) angles between any two pairs of agents are $\lambda_{\ell}$, where $\ell=\mathrm{EP},\mathrm{DE},$ and $\mathrm{DP}$. The engagement kinematics governing the relative motion between any two pairs can be expressed in polar coordinates as
@@ 2 Problem Formulation
$\dot{r}_{\mathrm{ED}}$ $=v_{\mathrm{D}}\cos\left(\gamma_{\mathrm{D}}-\lambda_{\mathrm{ED}}\right)-v_{\mathrm{E}}\cos\left(\gamma_{\mathrm{E}}-\lambda_{\mathrm{ED}}\right)$
$=v_{\mathrm{D}}\cos\delta_{\mathrm{DE}}-v_{\mathrm{E}}\cos\delta_{\mathrm{ED}}=v_{r_{\mathrm{ED}}},$ (2e)
$r_{\mathrm{ED}}\dot{\lambda}_{\mathrm{ED}}$ $=v_{\mathrm{D}}\sin\left(\gamma_{\mathrm{D}}-\lambda_{\mathrm{ED}}\right)-v_{\mathrm{E}}\sin\left(\gamma_{\mathrm{E}}-\lambda_{\mathrm{ED}}\right)$
$=v_{\mathrm{D}}\sin\delta_{\mathrm{DE}}-v_{\mathrm{E}}\sin\delta_{\mathrm{ED}}=v_{\lambda_{\mathrm{ED}}},$ (2f)
where $v_{r_{\mathrm{EP}}},v_{r_{\mathrm{DP}}},v_{r_{\mathrm{ED}}},v_{\lambda_{\mathrm{EP}}},v_{\lambda_{\mathrm{DP}}}$, and $v_{\lambda_{\mathrm{ED}}}$, represent the components of relative velocities of the relevant agents along and perpendicular to the corresponding LOS of their respective engagements. The quantities $\delta_{k}$ in (2) denote the corresponding lead angles and are defined as $\delta_{\mathrm{PE}}=\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}$, $\delta_{\mathrm{EP}}=\gamma_{\mathrm{E}}-\lambda_{\mathrm{EP}}$, $\delta_{\mathrm{PD}}=\gamma_{\mathrm{P}}-\lambda_{\mathrm{DP}}$, $\delta_{\mathrm{DP}}=\gamma_{\mathrm{D}}-\lambda_{\mathrm{DP}}$, $\delta_{\mathrm{ED}}=\gamma_{\mathrm{E}}-\lambda_{\mathrm{ED}}$, $\delta_{\mathrm{DE}}=\gamma_{\mathrm{D}}-\lambda_{\mathrm{ED}}$. Note that (2a)–(2b) describe the equations of motion between the evader and the pursuer, while (2c)–(2d) represent the same for the defender-pursuer pair. On the other hand, the cooperative engagement between the evader-defender pair is described by (2e)–(2f).
@@ 3 Main Results
[…]
@@
Consider the three-body engagement described using (2). If the defender maintains a positive closing speed with respect to the pursuer and attains an angle $\chi\in\left[{\pi}/{2},{3\pi}/{2}\right]$, then the pursuer’s capture is guaranteed before the evader could be captured.
@@
Referring to Figure 2, it is immediate that if $\chi\in\left({\pi}/{2},{\pi}\right)$ represents a scenario depicted in $\triangle$ P1E1D1. It follows that the side E1P1 is the longest. Hence, E1P1>>D1P1. If the evader-defender team cooperates such that the defender maintains a fixed $\chi\in\left({\pi}/{2},{\pi}\right)$ while also ensuring a positive closing speed with respect to the pursuer ($v_{r_{\mathrm{DP}}}<0$), the $\triangle$ P1E1D1 will shrink in proportion and may change its orientation, thereby generating smaller similar triangles. Eventually, D1P1 will degenerate to zero (or $r_{\mathrm{D}_{1}\mathrm{P}_{1}}\to 0$) while E1P1 (or $r_{\mathrm{E}_{1}\mathrm{P}_{1}}$) will still be positive. By a similar argument, it readily follows that $r_{\mathrm{D}_{2}\mathrm{P}_{2}}\to 0$ before $r_{\mathrm{E}_{2}\mathrm{P}_{2}}$ if $\chi\in\left({\pi},{3\pi}/{2}\right)$. The cases of $\chi={\pi}/{2},\pi$ are extremes but no different. When $\chi={\pi}/{2}$ (as in $\triangle$ P0E0D0), the side E0P0 is the hypotenuse and still the longest, so the same reasoning can be applied. Finally, when $\chi=\pi$, the defender is always directly between and on the pursuer-evader LOS, resulting in the pursuer’s capture by the defender prior to the interception of the evader by the pursuer. ∎
@@
$\chi=\pi+\lambda_{\mathrm{DP}}-\lambda_{\mathrm{ED}}$ (3)
for a general case. Toward this end, our goal is to ensure that the defender always attains a fixed angle, say, $\chi^{\star}\in\left[\pi/2,3\pi/2\right]$. Thus, the control objective is to nullify the error variable,
$\beta=\chi-\chi^{\star}=\pi+\lambda_{\mathrm{DP}}-\lambda_{\mathrm{ED}}-\chi^{\star}.$ (4)
Additionally, we also desire that $\beta\to 0$ within a time that is independent and uniform with respect to the initial three-body engagement geometry for a guaranteed pursuit evasion.
@@
The dynamics of $\beta$ has a relative degree of two with respect to the steering controls of the evader and defender.
@@
The derivative of $\beta$ can be obtained by differentiating (4) with respect to time and using the fact that $\dot{\pi}=\dot{\chi}^{\star}=0$ as
$\dot{\beta}=\dot{\lambda}_{\mathrm{DP}}-\dot{\lambda}_{\mathrm{ED}},$
which, on further differentiating with respect to time, yields
$\ddot{\beta}=\ddot{\lambda}_{\mathrm{DP}}-\ddot{\lambda}_{\mathrm{ED}}.$ (5)
Next, we obtain the values of $\ddot{\lambda}_{\mathrm{DP}}$ and $\ddot{\lambda}_{\mathrm{ED}}$. The value of $\ddot{\lambda}_{\mathrm{DP}}$ can be obtained by differentiating the LOS rate of evader-pursuer engagement given in (2d) with respect to time as
$r_{\mathrm{DP}}\ddot{\lambda}_{\mathrm{DP}}+\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}=v_{\mathrm{P}}\cos\delta_{\mathrm{PD}}\dot{\delta}_{\mathrm{PD}}-v_{\mathrm{D}}\cos\delta_{\mathrm{DP}}\dot{\delta}_{\mathrm{DP}}.$ (6)
Using the fact that $\dot{\delta}_{\mathrm{PD}}=\dot{\gamma}_{\mathrm{P}}-\dot{\lambda}_{\mathrm{DP}}$ and $\dot{\delta}_{\mathrm{DP}}=\dot{\gamma}_{\mathrm{D}}-\dot{\lambda}_{\mathrm{DP}}$, together with (1), the expression in (6) becomes
$r_{\mathrm{DP}}\ddot{\lambda}_{\mathrm{DP}}$ $=v_{\mathrm{P}}\cos\delta_{\mathrm{PD}}\left(\dfrac{a_{\mathrm{P}}}{v_{\mathrm{P}}}-\dot{\lambda}_{\mathrm{DP}}\right)-v_{\mathrm{D}}\cos\delta_{\mathrm{DP}}\left(\dfrac{a_{\mathrm{D}}}{v_{\mathrm{D}}}-\dot{\lambda}_{\mathrm{DP}}\right)$
$~-\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}.$ (7)